IBM just paid a $5 dividend and dividends are expected to grow at a perpetual rate of 8% for the foreseeable future. What is the value of the stock if investors require a 16% return?
Answer #1

Current price=D1/(Required return-Growth rate)

=(5*1.08)/(0.16-0.08)

which is equal to

=$67.50
Answer #2

Solution:

Given:

  • Dividend just paid (D₀): $5

  • Dividend growth rate (g): 8% (0.08)

  • Required return (r): 16% (0.16)

Objective:
Calculate the value of the stock using the Gordon Growth Model (Dividend Discount Model for perpetual growth).



Step 1: Calculate Next Year's Dividend (D₁)

Since dividends grow at 8%, the next dividend (D₁) is:

D1=D0×(1+g)=5×(1+0.08)=$5.40

Step 2: Apply the Gordon Growth Model

The stock value (P0) is given by:

P0=D1rg

Plug in the values:

P0=5.400.160.08=5.400.08=$67.50


Final Answer:

The value of the stock is 67.50.


answered by: anonymous